Is each ordered pair a solution of the given system of equations?
Show your work
2x + 5y=3
y= 7.5x; (1.5,0.2)

Answer by stanbon(75887)

Is each ordered pair a solution of the given system of equations?
Show your work
2x + 5y=3
y= 7.5x; (1.5,0.2)
-----
Substitute in each to see if you get a true statement.
2*1.5 + 5*0.2 = 3
3 + 1 = 3
False
--------------------
0.2 = 7.5*1.5
0.2 = 11.25
False
-----
(1.5,0.2) is not a solution of either equation.
=================================================
Cheers,
Stan H.
===============